If the sum of two numbers is 42. The second is 10 more than 3 times the first. Find the numbers.
a. 12,30
b. 8,34
c. 12,22
d. 10,32

Answer by rfer(16322)

x+y=42
y=3x+10
--------------
x+3x+10=42
4x=32
x=8
---------------
8+y=42
y=42-8
y=34
--------------
8+34=42